Laverny tosses an unbiased coin. He receives P100 if a head appears and he pays P40 it a tail appears Find the expected value and the variance of his gain.

1
Expert's answer
2022-03-11T06:07:04-0500

Let X= the amount of money which Laverny receives.


"\begin{matrix} x & 100 & -40 \\ p(x) & 0.5 & 0.5 \end{matrix}"


(a) Expected value "=E[X]"

"=0.5(100)+0.5(-40)=30"


(b) Variance

"E[X^2]=0.5(100)^2+0.5(-40)^2=5800"


"V(X)=E[X^2]-(E[X])^2=V(X)="

"=5800-(30)^2=4900"